OpenStudy (bballanic):

The message 29 34 39 57 45 65 2 3 17 22 54 74 26 39 29 4329 34 39 57 45 65 2 3 17 22 54 74 26 39 29 43 was encoded with matrix A. Decode the message.

OpenStudy (bballanic):

A = \[\left[\begin{matrix}1 & 1 \\ 2& 3\end{matrix}\right]\]

OpenStudy (bballanic):

A = \[\left[\begin{matrix}1 & 1 \\ 2& 3\end{matrix}\right]\]

OpenStudy (wholock):

Well, to start with, if the data set was encrypted with the a matrix, then to decode it you must use the inverse. to find the inverse, you must first find the determinant. Do you know how to do that?

OpenStudy (bballanic):

ad-bc ?

OpenStudy (bballanic):

I thought the inverse of A was \[\left[\begin{matrix}3 & -1 \\ -2 & 1\end{matrix}\right]\]

OpenStudy (bballanic):

I then multiplied the inverse of A with the message set in a 2 x 2 matrix and I got numbers higher than 26 (the alphabet was 26 letters) so I'm confused as to how to decode the numbers... how does 53 match up to a letter??? it doesn't... :/

OpenStudy (wholock):

Yes! That is correct! Now you just multiply the coded matrix by the inverse!

OpenStudy (bballanic):

has*

OpenStudy (bballanic):

I got numbers like 53, -24, 60, -21, 70, -25

OpenStudy (bballanic):

I feel like I went wrong somewhere I just don't know where

OpenStudy (wholock):

Just a minute, I have to do something, and I'll decode the matrix for you

OpenStudy (wholock):

Do you have about 20 minutes?

OpenStudy (bballanic):

no I have 8 mins before the timed quiz is up :/ but I guess I could just re-take it

OpenStudy (wholock):

Okay, just wait a min, I'll solve it right now!

OpenStudy (bballanic):

thanks so much! I know it takes a bit of time and is quite obnoxious so I appreciate that

OpenStudy (wholock):

Okay, um, could you clarify the coded matrix for me? how many rows and columns and the order pleas.

OpenStudy (bballanic):

29 39 45 2 17 54 26 29 34 57 65 3 22 74 39 43

OpenStudy (wholock):

Thank you!

OpenStudy (bballanic):

:)

OpenStudy (bballanic):

i have 2 and a half minutes D:

OpenStudy (wholock):

Okay 19 3 5 0 7 14 0 1 5 18 20 15 20 13 14

OpenStudy (wholock):

And that bottom on is 1 5 not 15